-15,-12,-9,-6,-3,......
solve the pattern,
please with steps.

Found 3 solutions by josgarithmetic, Alan3354, Theo:
Answer by josgarithmetic(39617)

Obvious arithmetic progression. The steps are trivial.

Answer by Alan3354(69443)

-15,-12,-9,-6,-3,......
solve the pattern,
please with steps.
-----------------------
Look at the difference between terms.

Answer by Theo(13342)

each time you're adding 3 to the previous number.
the formula is therefore An = An-1 + 3
A1 = -15
A2 = -15 + 3 = -12
A3 = -12 + 3 = -9
A4 = -9 + 3 = -6
A5 = -6 + 3 = -3
etc.....